# Graphs Meet AI Agents: Taxonomy, Progress, and Future Opportunities

### v1

@@ II-B Reinforcement Learning
Why RL is a suitable learning paradigm for AI agents? Firstly, reinforcement learning is inherently built on an ”interaction-feedback” loop, reflecting an agent’s continuous ”perceive-act” cycle. In this way, it can learn from experiences without relying heavily on labeled datasets. Secondly, RL pairs naturally with fast simulators, allowing agents to accumulate vast experience through self-play or exploration before deployment. Furthermore, RL encodes task objectives directly through reward design, letting the agent focus solely on maximizing return. This property enables the same learning framework to transfer flexibly across domains without the costly redesign of the objective function.
@@ II-C Large Language Models
In recent years, based on extensive world knowledge during large-scale pretraining, large language models (LLMs) have demonstrated great knowledge expression ability in natural language understanding and generation tasks [30]. They can generate coherent and logical text based on human prompts, performing various language tasks such as writing articles and answering questions.
@@
Planning (Sec. III), leaf-datasets, text width=9.6em [ Task Reasoning, leaf-datasets, text width=11em [ Knowledge Graph-Auxiliary
@@
Reasoning, leaf-datasets, text width=11.5em [ToT [38], RATT [39], GoT [40], Graph of Thoughts [41], RwG [42], modelnode-datasets, text width=38em] ] ] [ Task Decomposition, leaf-datasets, text width=11em [AgentKit [43], VillagerAgent [44], DynTaskMAS [45], Plan-over-Graph [46], DAG-Plan [47], LGC-MARL [48],
FGRL [49], GNN4TaskPlan [50], modelnode-datasets, text width=51.2em] ] [ Task Decision Searching, leaf-datasets, text width=11em [MCTS [51], M-MCTS [52], MENS-DT-RL [53], PromptAgent [54], LATS [55], MCGS [56], GBOP [57], CMCGS [58], modelnode-datasets, text width=51.2em] ] ] [ Graphs for Agent
Execution (Sec. IV), leaf-metrics,text width=9.6em [ Tool Usage, leaf-metrics, text width=11em [GRAPH-COT [59], GPTSwarm [16], ToolNet [17], Tool Fusion [60], ControlLLM [61], modelnode-metrics, text width=51.2em] ] [ Environment Interaction, leaf-metrics, text width=11em [ Heuristic-Based
@@
Memory (Sec. V), leaf-chall,text width=9.6em [ Memory Organization, leaf-chall, text width=11em [AriGraph [74], IKG [75], Graphusion [76], MemGraph [77], Mind Map [78], StructuralMemory [79], DAMCS [80],
GraphRAG [81], KG-Retriever [82], modelnode-chall, text width=51.2em] ] [ Memory Retrieval, leaf-chall, text width=11em [G-Retriever [83], GFM-RAG [84], SubgraphRAG [85], LightRAG [86], GRAG [87], PathRAG [88], modelnode-chall, text width=51.2em] ] [ Memory Maintenance, leaf-chall, text width=11em [A-MEM [89], AriGraph [74], DAMCS [80], Zep [90], HippoRAG [91], LightRAG [86], KG-Agent [92], InstructRAG [93], modelnode-chall, text width=51.2em] ] ] [ Graphs for Multi-Agent
@@
Passing, leaf-methods, text width=11em [ Task-Specific
@@
Optimization, leaf-methods, text width=11em [DICG [105], DiscoGraph [106], CoPU [107], DMCG [108], G2ANet [109], MAGIC [110], MAGE-X [111],
CommFormer [112], G-Designer [113], GNN-VAE [114], DyLAN [115], HGRL [116], GPTSwarm [16], modelnode-methods, text width=51.2em] ] ] [ Agents for Graph
@@
Synthesis, leaf-app, text width=11em [GPA [117], BIGENE [118], LLM4GraphGen [119], GAG [120], IGDA [121], Plan-over-Graph [46], GraphMaster [122], modelnode-app, text width=51.2em] ] [ Graph Understanding, leaf-app, text width=11em [Policy-GNN [123], SUGAR [124], MAG-GNN [125], AgentNet [126], GAgN [127], GL-Agent [128],
GraphGPT [129], GA [130], LLaGA [131], GraphInstruct [132], GraphWiz [133], GraphAgent-Reasoner [134],
GraphAgent [135], GraphTeam [136], modelnode-app, text width=51.2em] ] ] ]
@@ III-C Task Decision Searching
Task decision searching is an important process in the planning stage of AI agents, involving the making of sequential decisions in complex environments to achieve specific goals [144]. Search algorithms involve state transitions within the decision space, which naturally form a graph structure with transitions between states. These states and their transitions form a state space graph (SSG). Formally, each node in the SSG represents a state, with the node’s properties being the state’s information, such as its textual description or parameters. The transitions between states serve as the edges between nodes. Note that the edge relationships in the SSG are typically optimized by the search algorithm itself, rather than being predefined. It involves an agent’s process of exploring different states and actions to dynamically generate the SSG within an environment, based on the current state and environment, to identify a sequence of actions that maximizes long-term rewards.
@@ VI-B Coordination Topology Optimization
Edge Importance Measurement. In an overall pipeline, these methods first initialize a fully connected graph between multiple agents, such as ACG, and then learn potential edges via attention mechanisms/importance scores or learnable parameterized edge weights, like DICG [105], DiscoGraph [106], CoPU [107], and DMCG [108]. Furthermore, as fully connected graphs are computationally inefficient due to the high topology density, some methods convert them back to discrete graphs by threshold binarization (edges exist if weights exceed a threshold, otherwise not) or differentiable Gumbel-Softmax discretization [160], such as G2ANet [109], MAGIC [110], MAGE-X [111], and CommFormer [112].
Graph Auto-Encoder Optimization. Graph auto-encoders (GAEs) encode node representations via an encoder layer and optimize specific tasks via a decoder layer. Many graph learning works use the GAE architecture for self-supervised link prediction between nodes [161, 162]. The core idea is to learn node representations with the encoder and to predict the links with a decoder. Based on this insight, GAE can be used for agent communication topology learning by predicting the high-confidence edges that exist between agents. Related works like G-Designer [113] and GNN-VAE [114], through graph variational autoencoders [163], robustly predict whether edges exist between agent nodes to determine the optimal multi-agent coordination topology.
@@ VII-B Graph Understanding
LLM Agents. With the powerful knowledge of LLMs, many works are exploring how to design and tune LLM agents to fully leverage the capabilities of LLMs in graph-modeling tasks. The core of these works is to post-train LLMs on graph data. Thus, LLM agents can automatically understand graph data through text-described graph information and carry out downstream tasks like node classification. GL-Agent [128] and GraphGPT [129] initially propose LLM frameworks for agentic graph learning on various graphs. GA [130] uses LLM as the foundation model of an agent, and employs a memory mechanism to store nodes or edges embeddings for retrieval augmentation. LLaGA [131] further transforms graph inputs into a token space comprehensible to LLMs. This transformation allows the use of the inherent reasoning abilities of LLMs for graph-related tasks without modifying the LLM parameters. GraphInstruct [132] constructs diverse graph generation processes and descriptions, and proposes a step-by-step masking training strategy for LLM supervised fine-tuning. Besides tuning to make models generate reasoning paths based on graph structure and text prompts, GraphWiz [133] further improves reasoning ability by using DPO [166] to compare reasoning paths.
@@ IX-D Multimodal Agents
The agent paradigm using LLMs as foundation models has achieved remarkable performance in executing tasks in the language space. Research on agents and LLMs is still constantly evolving. For example, there have been some breakthroughs in large vision language models [231] or multimodal large language models [232] recently. They can also serve as the foundation model for next-generation AI agents. Multimodal agents will surpass current systems centered on language or single modalities, evolving into unified entities capable of understanding and integrating information streams such as text, vision, and speech [233, 234]. Such agents should make decisions and learn autonomously from heterogeneous data sources. In this evolution, graph learning can play a vital role as multimodal data are inherently interlinked with complex and dynamic relationships. Graph techniques can abstract and connect different modalities, seamlessly integrating modality entities from the external world, perceived objects, and environmental states [235, 11].

### v2

@@ II-B Reinforcement Learning
Why RL is a suitable learning paradigm for AI agents? Firstly, reinforcement learning is inherently built on an ”interaction-feedback” loop, reflecting an agent’s continuous ”perceive-act” cycle. In this way, it can learn from experiences without relying heavily on labeled datasets. Secondly, RL pairs naturally with fast simulators, allowing agents to accumulate vast experience through self-play or exploration before deployment. Furthermore, RL encodes task objectives directly through reward design, letting the agent focus solely on maximizing return. These enable RL agents to flexibly design customized learning strategies according to the characteristics of specific tasks and environments.
@@ II-C Large Language Models
In recent years, based on extensive world knowledge during large-scale pretraining, large language models (LLMs) have demonstrated great knowledge expression ability in natural language understanding and generation tasks [30]. Although RL can flexibly shape agents for different scenarios through reward-based learning, RL agents need to be trained individually for each specific task. LLM can serve as an agent’s foundation model for various tasks without costly re-training, supported by the pre-trained foundational knowledge.
@@
Planning (Sec. III), leaf-datasets, text width=9.6em [ Task Reasoning, leaf-datasets, text width=10.5em [ Knowledge Graph-Auxiliary
@@
Reasoning, leaf-datasets, text width=11.5em [ToT [38], RATT [39], GoT [40], Graph of Thoughts [41], RwG [42], modelnode-datasets, text width=38em] ] ] [ Task Decomposition, leaf-datasets, text width=10.5em [AgentKit [43], VillagerAgent [44], DynTaskMAS [45], Plan-over-Graph [46], DAG-Plan [47], LGC-MARL [48],
FGRL [49], GNN4TaskPlan [50], modelnode-datasets, text width=51.2em] ] [ Task Decision Searching, leaf-datasets, text width=10.5em [MCTS [51], M-MCTS [52], MENS-DT-RL [53], PromptAgent [54], LATS [55], MCGS [56], GBOP [57], CMCGS [58], modelnode-datasets, text width=51.2em] ] ] [ Graphs for Agent
Execution (Sec. IV), leaf-metrics,text width=9.6em [ Tool Usage, leaf-metrics, text width=10.5em [GRAPH-COT [59], GPTSwarm [16], ToolNet [17], Tool Fusion [60], ControlLLM [61], modelnode-metrics, text width=51.2em] ] [ Environment Interaction, leaf-metrics, text width=10.5em [ Heuristic-Based
@@
Memory (Sec. V), leaf-chall,text width=9.6em [ Memory Organization, leaf-chall, text width=10.5em [AriGraph [74], IKG [75], Graphusion [76], MemGraph [77], Mind Map [78], StructuralMemory [79], DAMCS [80],
GraphRAG [81], KG-Retriever [82], modelnode-chall, text width=51.2em] ] [ Memory Retrieval, leaf-chall, text width=10.5em [G-Retriever [83], GFM-RAG [84], SubgraphRAG [85], LightRAG [86], GRAG [87], PathRAG [88], modelnode-chall, text width=51.2em] ] [ Memory Maintenance, leaf-chall, text width=10.5em [A-MEM [89], AriGraph [74], DAMCS [80], Zep [90], HippoRAG [91], LightRAG [86], KG-Agent [92], InstructRAG [93], modelnode-chall, text width=51.2em] ] ] [ Graphs for Multi-Agent
@@
Passing, leaf-methods, text width=10.5em [ Task-Specific
@@
Optimization, leaf-methods, text width=10.5em [DICG [105], DiscoGraph [106], CoPU [107], DMCG [108], G2ANet [109], MAGIC [110], MAGE-X [111],
CommFormer [112], AgentPrune [113], AGP [114], G-Designer [115], GNN-VAE [116], DyLAN [117],
HGRL [118], GPTSwarm [16], modelnode-methods, text width=51.2em] ] ] [ Agents for Graph
@@
Synthesis, leaf-app, text width=10.5em [GPA [119], BIGENE [120], LLM4GraphGen [121], GAG [122], IGDA [123], Plan-over-Graph [46], GraphMaster [124], modelnode-app, text width=51.2em] ] [ Graph Understanding, leaf-app, text width=10.5em [Policy-GNN [125], SUGAR [126], MAG-GNN [127], AgentNet [128], GAgN [129], GL-Agent [130],
GraphGPT [131], GA [132], OFA [133], ZeroG [134], LLaGA [135], GraphInstruct [136], GraphWiz [137],
GraphAgent-Reasoner [138], GraphAgent [139], GraphTeam [140], modelnode-app, text width=51.2em] ] ] ]
@@ III-C Task Decision Searching
Task decision searching is an important process in the planning stage of AI agents, involving the making of sequential decisions in complex environments to achieve specific goals [148]. Search algorithms involve state transitions within the decision space, which naturally form a graph structure with transitions between states. These states and their transitions form a state space graph (SSG). Formally, each node in the SSG represents a state, with the node’s properties being the state’s information, such as its textual description or parameters. The transitions between states serve as the edges between nodes. Note that the edge relationships in the SSG are typically optimized by the search algorithm itself, rather than being predefined. It involves an agent’s process of exploring different states and actions to dynamically generate and update the SSG within an environment, based on the current state and environment, to identify a sequence of actions that maximizes long-term rewards.
@@ VI-B Coordination Topology Optimization
Edge Importance Measurement. In an overall pipeline, these methods first initialize a fully connected or rule-based topology between multiple agents. Then, on the one hand, many works select helpful edges via attention mechanisms, importance scores, or learnable parameterized edge weights, like DICG [105], DiscoGraph [106], CoPU [107], and DMCG [108]. Building on this line, as parameterized connected graphs are computationally inefficient due to the high topology density, some methods convert them back to discrete graphs by threshold binarization (edges exist if weights exceed a threshold, otherwise not) or differentiable Gumbel-Softmax discretization [164], such as G2ANet [109], MAGIC [110], MAGE-X [111], and CommFormer [112]. On the other hand, some works directly utilize trainable graph masking for redundant and unnecessary communication node or edge pruning. For example, AgentPrune [113] identifies key multi-agent connections through low-rank-principled graph masking, and AGP [114] dynamically determines which agent nodes participate in specific tasks through a node masking head.
Graph Auto-Encoder Optimization. Graph auto-encoders (GAEs) encode node representations via an encoder layer and optimize specific tasks via a decoder layer. Many graph learning works use the GAE architecture for self-supervised link prediction between nodes [165, 166]. The core idea is to learn node representations with the encoder and to predict the links with a decoder. Based on this insight, GAE can be used for agent communication topology learning by predicting the high-confidence edges that exist between agents. Related works like G-Designer [115] and GNN-VAE [116], through graph variational autoencoders [167], robustly predict whether edges exist between agent nodes to determine the optimal multi-agent coordination topology. AGP [114] employs GAE for soft-pruning on sampled communication topologies and supplying supervision signals for the subsequent optimization within the framework.
@@ VII-B Graph Understanding
LLM Agents. With the powerful knowledge of LLMs, many works are exploring how to design and tune LLM agents to fully leverage the capabilities of LLMs in graph-modeling tasks. The core of these works is to post-train LLMs on graph data. Thus, LLM agents can automatically understand graph data through text-described graph information and carry out downstream tasks like node classification. GL-Agent [130] and GraphGPT [131] initially propose LLM frameworks for agentic graph learning on various graphs. GA [132] uses LLM as the foundation model of an agent, and employs a memory mechanism to store nodes or edges embeddings for retrieval augmentation. Additionally, OFA [133] and ZeroG [134] use the LLM agent to map nodes from different graphs into a unified semantic space to ensure the generalization ability across datasets and graph understanding tasks. LLaGA [135] further transforms graph inputs into a token space comprehensible to LLMs. This transformation allows the use of the inherent reasoning abilities of LLMs for graph-related tasks without modifying the LLM parameters. GraphInstruct [136] constructs diverse graph generation processes and descriptions, and proposes a step-by-step masking training strategy for LLM supervised fine-tuning. Besides tuning to make models generate reasoning paths based on graph structure and text prompts, GraphWiz [137] further improves reasoning ability by using DPO [170] to compare reasoning paths.
@@ IX-D Multimodal Agents
The agent paradigm using LLMs as foundation models has achieved remarkable performance in executing tasks in the language space. Research on agents and LLMs is still constantly evolving. For example, there have been some breakthroughs in large vision language models [235] or multimodal large language models [236] recently. They can also serve as the foundation model for next-generation AI agents. Multimodal agents will surpass current systems centered on language or single modalities, evolving into unified entities capable of understanding and integrating information streams such as text, vision, and speech [237, 238]. Such agents should make decisions and learn autonomously from heterogeneous data sources. In this evolution, graph learning can play a vital role as multimodal data are inherently interlinked with complex and dynamic relationships. Graph techniques can abstract and connect different modalities, seamlessly integrating and aligning modality entities from the external world, perceived objects, and environmental states [239, 11, 240].